OpenStudy (anonymous):

A line has a rise of 3 and a run of 12. What is the slope?

OpenStudy (quietus):

same method

OpenStudy (bahrom7893):

Slope = rise/run

OpenStudy (quietus):

please do try them on your own, after knowing the method ^

OpenStudy (imstuck):

I recommend setting this up as a ratio. Like this:

OpenStudy (imstuck):

\[\frac{ 3 }{ 12 }=\frac{ ? }{ ? }\]This one can just be reduced. The simplest form of the answer is your slope.